# Deep Joint CSI Estimation-Feedback-Precoding for MU-MIMO OFDM Systems

### v1

@@ Abstract
As the number of antennas in frequency-division duplex (FDD) multiple-input multiple-output (MIMO) systems increases, acquiring channel state information (CSI) becomes increasingly challenging due to limited spectral resources and feedback overhead. In this paper, we propose an end-to-end network that conducts joint design with pilot design, CSI estimation, CSI feedback, and precoding design in the multi-user MIMO orthogonal frequency-division multiplexing (OFDM) scenario. Multiple communication modules are jointly designed and trained with a common optimization objective to prevent mismatches between modules and discrepancies between individual module objectives and the final system goal. Experimental results demonstrate that, under the same feedback and CE overheads, the proposed joint multi-module end-to-end network achieves a higher multi-user downlink spectral efficiency than traditional algorithms based on separate architecture and partially separated artificial intelligence-based network architectures under comparable channel quality. Furthermore, compared to conventional separate architecture, the proposed network architecture with joint architecture reduces the computational burden and model storage overhead at the UE side, facilitating the deployment of low-overhead multi-module joint architectures in practice. While slightly increasing storage requirements at the base station, it reduces computational complexity and precoding design delay, effectively reducing the effects of channel aging challenges.
Index Terms:
CSI feedback, precoding design, CSI estimation, deep joint source-channel coding.
@@ I Introduction
With the significant performance improvements brought by massive multiple-input multiple-output (MIMO) technology in 5G systems, and the growing demands for higher spectrum utilization, increased communication throughput, and faster transmission rates in the vision of 6G [1], MIMO technology and its evolving variants are expected to become key technologies for 6G. The development of MIMO technology, such as extremely large-scale MIMO [2] and reconfigurable intelligent surface [3] technologies, offers valuable insights into the application of MIMO in 6G systems. Additionally, research into terahertz band communication [4] provides theoretical support for MIMO systems that can accommodate more antennas and more effectively exploit spatial resources. However, to fully harness the potential of MIMO technology, obtaining channel state information (CSI) with low overhead is a critical challenge. In time-division duplex (TDD) systems, channel reciprocity between uplink and downlink channels operating in the same frequency band simplifies downlink CSI acquisition. Conversely, in frequency-division duplex (FDD) systems, where channel reciprocity is weak, the increased number of antennas presents significant challenges for downlink CSI estimation and feedback with low overhead and low latency.
@@
Traditional methods for CSI feedback include the codebook-based feedback method, commonly applied in 5G, and the compressed sensing (CS) method based on sparsity assumptions [5]. However, the codebook-based feedback approach faces scalability challenges as the number of antennas increases. Ensuring the same reconstruction accuracy requires a larger codebook size, leading to increased complexity in codeword search and higher feedback overhead. Similarly, the CS-based CSI feedback method is constrained by the sparsity assumption, making it difficult to achieve theoretical performance. Additionally, the complex reconstruction algorithms in CS methods introduce significant computational overhead and delay, making them unsuitable for fast time-varying channels prone to CSI aging.
@@
With the promising performance of artificial intelligence (AI) techniques in fields such as image compression and transmission [6], and signal processing [7], AI-based methods are increasingly adopted for communication tasks, particularly CSI feedback [8, 9]. These methods treat the CSI as an image, where the real and imaginary components are considered as separate channels. AI-based techniques exploit the correlation of CSI in the frequency and spatial domains to compress high-dimensional CSI into low-dimensional feature vectors using an auto-encoder at the transmitter. The compressed features are fed back, and the CSI is reconstructed at the receiver using a corresponding decoder. This CSI feedback process is considered an image compression and reconstruction task, trained end-to-end from user equipment (UE) to the base station (BS), with reconstruction accuracy as the optimization objective to update trainable parameters. By replacing traditional source coding and decoding modules with AI-based compression and reconstruction modules, various network architectures have been proposed to achieve low-overhead, high-accuracy CSI feedback, addressing key limitations of traditional methods [10, 11, 12].
@@
The use of AI networks to replace traditional source coding and decoding modules in communication architectures is referred to as a separate source-channel coding (SSCC) architecture. In this approach, the design of the separate architecture allows the network to be trained without considering the impact of channel coding and decoding on performance, which results in the network performance being highly affected by channel quality. In contrast, the authors of [13] proposed a deep joint source-channel coding (DJSCC) architecture for CSI feedback, which jointly considers source coding, channel coding, channel decoding, and source decoding. This approach abandons the traditionally modular design seen in communication systems. By leveraging the powerful nonlinear capabilities of AI, the DJSCC module provides more stable and reliable CSI reconstruction performance compared to the SSCC architecture.
@@
The challenges encountered by separate architectures also arise in both the upstream task (CSI estimation) and the downstream task (precoding design) of CSI feedback. Inaccurate CSI estimation can cause a mismatch in the input distribution of the CSI compression module between the training and testing phases, leading to inaccurate CSI feedback. In our previous work [14], we explored several joint multi-module end-to-end networks for the CSI feedback task, demonstrating the benefits of joint multi-module design. In [10, 11, 15], AI-based two-sided CSI compression and reconstruction modules are designed and jointly trained to perform the CSI feedback task, typically resulting in improved CSI reconstruction performance compared to one-sided feedback architectures in [8, 16, 17] and codebook-based feedback architectures in traditional algorithms. In [13], the CSI compression and reconstruction module is jointly designed with the channel coding and decoding module to mitigate the “cliff effect” and reduce retransmission delays. In our previous work [18], we proposed a joint feedback and precoding network (JFPNet) for multi-user (MU) MIMO scenarios that simultaneously addresses both CSI feedback and its downstream task, precoding design. The source coding and channel coding modules are jointly designed on the UE side, while the channel decoding, source decoding, and precoding design modules for MU scenarios are jointly designed on the BS side. According to the experimental results in [18], this joint multi-module architecture outperforms the separate architecture for the CSI feedback and precoding design task. In [19], an end-to-end multi-module joint architecture was designed to simultaneously address upstream and downstream tasks, including pilot length design, CE, CSI feedback, and beamforming precoding. However, its pilot sequences are randomly generated from a complex unit sphere, limiting the ability to fully utilize channel information and exploit the learning capacity of the end-to-end architecture. Additionally, its downlink CE process relies on artificially introduced noise to simulate CE errors, which are inconsistent with the actual errors caused by specific CE methods, resulting in performance degradation. Both the joint CSI feedback and estimation architecture in [20] and the joint CSI estimation, feedback, and precoding design architecture in [21, 22] address the end-to-end performance loss caused by mismatches between the distribution of the estimated and trained CSI. However, both overlook the potential “cliff effect” caused by the omission of channel coding and decoding, as well as the added retransmission overhead and delay when channel quality degrades.
To address this gap, we propose a joint CSI estimation, feedback, and precoding design network (JEFPNet). In summary, our contributions to this work are as follows:
@@
    In a MU MIMO-OFDM system, we propose an end-to-end network that jointly addresses CSI feedback, its upstream task (CSI estimation), and its downstream task (precoding design). The proposed framework considers the entire closed-loop system, starting from the pilot design at the BS, to the CSI estimation and joint source-channel coding at the UE, and finally to the joint source-channel decoding and precoding design at the BS. Only one joint multi-module designed network is deployed at each device terminal.
@@
    By jointly designing the pilot, CSI estimation, and CSI compression modules, the framework extracts task-relevant semantic information for transmission through end-to-end training, which indirectly reduces the source compression rate and improves feedback efficiency. Additionally, by avoiding explicit CE, the proposed JEFPNet decreases both storage and computational overhead at the UE.
@@
    Performance comparisons between joint and separate architectures are conducted through simulation experiments. The results demonstrate the performance gains and stability of the joint architecture. The proposed JEFPNet proves to be more adaptive than separate architectures, particularly in the presence of large CSI estimation and feedback errors and is capable of designing more efficient precoding vectors than separate architecture.
@@
[…]
@@ II-A System Model
In this paper, we consider a MIMO OFDM system with $K$ users, where each user is equipped with a single omnidirectional antenna, and the BS employs an array antenna with $N_{t}$ elements. The system operates with $N_{c}$ subcarriers. In this system, we simultaneously consider downlink CSI estimation, downlink CSI feedback in the uplink channel, and downlink precoding design. The downlink CSI of the $k$-th user can be expressed as $\mathbf{H}_{d}(k)\in\mathbb{C}^{N_{c}\times N_{t}}$, and the corresponding uplink feedback channel can be expressed as $\mathbf{H}_{u}(k)\in\mathbb{C}^{N_{c}\times N_{t}}$.
For the downlink CSI estimation phase, we consider the pilot-assisted CE method, that downlink CSI is estimated at the UE with the help of the known pilot sequence for transmitter and receiver. For the pilot sequence, in the time domain, we use sequential $L$ OFDM symbols for CE. To reduce pilot overhead, pilots are placed at equal $g$ intervals in the frequency domain, while the non-pilot positions are utilized for data or other channel control information transmission. $M$ subcarriers are sampled at equal intervals $g$ from the $N_{c}$ subcarriers to place the pilots, and the indices of these subcarriers form the set of pilot sequences, denoted as =$\mathcal{M}=\{\mathcal{M}_{1},\cdots,\mathcal{M}_{M}\}\in\mathbb{R}$, where $M=\lfloor{N_{c}/g}\rfloor$. In this paper, user interference during feedback is not considered. Consequently, the same pilot symbols are employed across different users. The pilot sequence at the $\mathcal{M}_{m}$-th subcarrier can be expressed as $\mathbf{P}_{\mathcal{M}_{m}}=\left[{\mathbf{p}_{\mathcal{M}_{m}}^{(1)},\cdots,\mathbf{p}_{\mathcal{M}_{m}}^{(L)}}\right]\in\mathbb{C}^{N_{t}\times L}$.
@@
At the BS, we use an maximal ratio combining (MRC) receiver for detecting signals. The combining matrix is derived by the estimated uplink CSI. In this paper, we assume perfect uplink CE. Therefore, the combining matrix of the $z$-th subcarrier $\mathbf{W}_{z}(k)$ can be expressed as

### v2

# Joint CSI Estimation-Feedback-Precoding via DJSCC for MU-MIMO OFDM Systems

@@ Abstract
As the number of antennas in frequency-division duplex (FDD) multiple-input multiple-output (MIMO) systems increases, acquiring channel state information (CSI) becomes increasingly challenging due to limited spectral resources and feedback overhead. In this paper, we investigate the impact of the feedback channel on CSI feedback in a multi-user MIMO orthogonal frequency-division multiplexing (OFDM) scenario, where the received downlink pilot signal is directly utilized as the source for CSI feedback in a joint design with CSI feedback and precoding. Considering the influence of the feedback channel, we propose an end-to-end joint CSI estimation–feedback–precoding network based on a deep joint source–channel coding architecture with an adaptive number of users. Experimental results demonstrate that, under the same feedback and CSI estimation overheads, the proposed joint multi-module end-to-end network achieves a higher multi-user downlink spectral efficiency than traditional algorithms based on separate architecture and partially separated artificial intelligence-based network architectures under comparable channel quality. Furthermore, compared to conventional separate architecture, the proposed network architecture with joint architecture reduces the computational burden and model storage overhead at the UE side, facilitating the deployment of low-overhead multi-module joint architectures in practice. Meanwhile, the network designed at the BS achieves user-number adaptability without increasing the number of trainable parameters, thereby reducing both model storage and distribution overhead by requiring only a single set of parameters for different numbers of users. While slightly increasing storage requirements at the base station, it reduces computational complexity and precoding design delay, effectively reducing the effects of channel aging challenges.
@@ I Introduction
With the significant performance improvements brought by massive multiple-input multiple-output (MIMO) technology in 5G systems, and the growing demands for higher spectrum utilization, increased communication throughput, and faster transmission rates in the vision of 6G [Survey_MIMO], MIMO technology and its evolving variants are expected to become key technologies for 6G. The development of MIMO technology, such as extremely large-scale MIMO [XL_MIMO] and reconfigurable intelligent surface [RIS] technologies, offers valuable insights into the application of MIMO in 6G systems. Additionally, research into terahertz band communication [THz] provides theoretical support for MIMO systems that can accommodate more antennas and more effectively exploit spatial resources. However, to fully harness the potential of MIMO technology, obtaining channel state information (CSI) with low overhead is a critical challenge. In time-division duplex (TDD) systems, channel reciprocity between uplink and downlink channels operating in the same frequency band simplifies downlink CSI acquisition. Conversely, in frequency-division duplex (FDD) systems, where channel reciprocity is weak, the increased number of antennas presents significant challenges for downlink CSI estimation (CE) and feedback with low overhead and low latency.
@@
Traditional methods for CSI feedback include the codebook-based feedback method, commonly applied in 5G, and the compressed sensing (CS) method based on sparsity assumptions [Sparse_Assume]. However, the codebook-based feedback approach faces scalability challenges as the number of antennas increases. Ensuring the same reconstruction accuracy requires a larger codebook size, leading to increased complexity in codeword search and higher feedback overhead. Similarly, the CS-based CSI feedback method is constrained by the sparsity assumption, making it difficult to achieve theoretical performance. Additionally, the complex reconstruction algorithms in CS methods introduce significant computational overhead and delay, making them unsuitable for fast time-varying channels prone to CSI aging.
@@
With the promising performance of artificial intelligence (AI) techniques in fields such as image compression and transmission [AI_Image_Process], and signal processing [AI_Signal_Process], AI-based methods are increasingly adopted for communication tasks, particularly CSI feedback [CsiNet, CSI_FB_Overview, One_Bit_Superimposed_CSI_FB, LOS_Superimposed_CSI_FB, GAT_Superimposed_CSI_FB]. In [One_Bit_Superimposed_CSI_FB], the feedback overhead can be reduced by superimposing downlink CSI onto the uplink signal; however, this approach introduces a new challenge, i.e., interference between the superimposed CSI feedback and the original uplink signal. To mitigate this challenge, [One_Bit_Superimposed_CSI_FB] proposed an AI-based fusion learning scheme using 1-bit CS, which leverages the partial reciprocity of uplink and downlink channels for CSI feedback and reconstruction. Another line of research retains the resources allocated for CSI feedback but seeks to enable the extraction of as many low-dimensional CSI features as possible without compromising reconstruction performance through advanced network design or the incorporation of auxiliary information, balancing low overhead and high reconstruction accuracy. These methods treat the CSI as an image, where the real and imaginary components are considered as separate channels. AI-based techniques exploit the correlation of CSI in the frequency and spatial domains to compress high-dimensional CSI into low-dimensional feature vectors using an auto-encoder at the transmitter. The compressed features are fed back, and the CSI is reconstructed at the receiver using a corresponding decoder. This CSI feedback process is considered an image compression and reconstruction task, trained end-to-end from user equipment (UE) to the base station (BS), with reconstruction accuracy as the optimization objective to update trainable parameters. By replacing traditional source coding and decoding modules with AI-based compression and reconstruction modules, various network architectures have been proposed to achieve low-overhead, high-accuracy CSI feedback, addressing key limitations of traditional methods [CsiNet_plus, SwinTransformer, Transformer_CSI].
@@
The use of AI networks to replace traditional source coding and decoding modules in communication architectures is referred to as a separate source-channel coding (SSCC) architecture. In this approach, the design of the separate architecture allows the network to be trained without considering the impact of channel coding and decoding on performance, which results in the network performance being highly affected by channel quality. In contrast, the authors of [ADJSCC] proposed a deep joint source-channel coding (DJSCC) architecture for CSI feedback, which jointly considers source coding, channel coding, channel decoding, and source decoding. This approach abandons the traditionally modular design seen in communication systems. By leveraging the powerful nonlinear capabilities of AI, the DJSCC module provides more stable and reliable CSI reconstruction performance compared to the SSCC architecture.
TABLE I: Comparison of Existing Studies and the Proposed Network Architecture
Ref.
Pilot
Design
Channel
Estimation
CSI
Compression
Channel
Coding
Channel
Decoding
CSI
Construction
Precoding
Design
Feedback
Method
[CsiNet_plus] Quantification + Ideal Feedback
[ADJSCC] DJSCC
[JFPNet] DJSCC
[Joint_Pilot_CE_FB_BF] Ideal Feedback
[CEFNet] Quantification + Ideal Feedback
[ICC_JEFPNet] Quantification + Ideal Feedback
[YuWei_JEFPNet] Quantification + Ideal Feedback
[Joint_PD_FB_BF] Quantification + Ideal Feedback
[MU_Adaptive] Quantification + Ideal Feedback
Proposed DJSCC
@@
The challenges encountered by separate architectures also arise in both the upstream task (CE) and the downstream task (precoding design) of CSI feedback. Inaccurate CE can cause a mismatch in the input distribution of the CSI compression module between the training and testing phases, leading to inaccurate CSI feedback. In [CsiNet_plus, Implicit_CSI_FB], AI-based two-sided CSI compression and reconstruction modules are designed and jointly trained to perform the CSI feedback task, typically resulting in improved CSI reconstruction performance compared to one-sided feedback architectures in [CsiNet, CFnet] and codebook-based feedback architectures in traditional algorithms. In [ADJSCC], the CSI compression and reconstruction module is jointly designed with the channel coding and decoding module to mitigate the “cliff effect” and reduce retransmission delays. In our previous work [JFPNet], we proposed a joint feedback and precoding network (JFPNet) for multi-user (MU) MIMO scenarios that simultaneously addresses both CSI feedback and its downstream task, precoding design. The source coding and channel coding modules are jointly designed on the UE side, while the channel decoding, source decoding, and precoding design modules for MU scenarios are jointly designed on the BS side. According to the experimental results in [JFPNet], this joint multi-module architecture outperforms the separate architecture for the CSI feedback and precoding design task. However, this work assumes that acquiring downlink CSI at the UE is ideal, and that multiple networks must be trained and stored to accommodate different numbers of users. In [Joint_Pilot_CE_FB_BF], an end-to-end multi-module joint architecture was designed to simultaneously address upstream and downstream tasks, including pilot length design, CE, CSI feedback, and beamforming precoding. However, its pilot sequences are randomly generated from a complex unit sphere, limiting the ability to fully utilize channel information and exploit the learning capacity of the end-to-end architecture. Additionally, its downlink CE process relies on artificially introduced noise to simulate CE errors, which are inconsistent with the actual errors caused by specific CE methods, resulting in performance degradation. Both the joint CSI feedback and estimation architecture in [CEFNet] and the joint CE, feedback, and precoding design architecture in [ICC_JEFPNet, YuWei_JEFPNet, Joint_PD_FB_BF, MU_Adaptive] address the end-to-end performance loss caused by mismatches between the distribution of the estimated and trained CSI. However, both overlook the potential “cliff effect” caused by the omission of channel coding and decoding , as shown in Table. I, as well as the added retransmission overhead and delay when channel quality degrades.
The impact of the “cliff effect” in CSI feedback on system performance has been thoroughly evaluated in [ADJSCC], and the proposed DJSCC architecture offers a new solution paradigm to mitigate this problem [JFPNet, DJSCC_FB_1, DJSCC_FB_2]. However, the system architecture remains to be explored when the source changes from the ideally acquired CSI, as considered in [ADJSCC, JFPNet, DJSCC_FB_1], to the more realistic non-ideal CSI obtained through CE or directly from the received pilot signal. Particularly in communication scenarios with a variable number of users, designing a bilateral framework for CE, CSI feedback, and precoding that is both user-number-adaptive and deployment-friendly is crucial to advancing the practical application of the DJSCC architecture. To address this challenge, and with channel coding taken into account, we propose a joint design of CE, CSI feedback, and precoding with user-number-adaptive capabilities. In summary, our contributions to this work are as follows:
@@
    In a MU MIMO orthogonal frequency-division multiplexing (OFDM) system, under consideration of the impact of channel coding on network performance, we propose an DJSCC-based end-to-end network that jointly addresses CSI feedback, its upstream task (CE), and its downstream task (precoding design). The proposed framework considers the entire closed-loop system, starting from the pilot design at the BS, to the CE and joint source-channel coding at the UE, and finally to the joint source-channel decoding and precoding design at the BS. Only one joint multi-module designed network is deployed at each device terminal.
@@
    Considering a variable number of users, we design a joint CE, CSI feedback, and precoding network based on a Transformer architecture [Transformer] that is adaptive to the number of UE. The number of stored parameters in the proposed network remains constant regardless of the number of users, and within a defined user range, its performance approximates that of networks trained with a fixed number of users, eliminating the need to switch models. This design significantly reduces model storage overhead, switching delays and model distribution overhead form the BS to the UE when the number of users changes.
@@
    Performance comparisons between joint and separate architectures with DJSCC architecture under different numbers of UEs are conducted through simulation experiments. By jointly designing the pilot, CE, and CSI compression modules, the proposed framework extracts task-oriented semantic information for transmission via end-to-end training, thereby enhancing feedback efficiency. The simulation results demonstrate both the performance gains and the stability of the joint architecture. In particular, the proposed JEFPNet exhibits stronger adaptability than separate architectures, especially in the presence of significant CE and feedback errors, and is capable of designing more efficient precoding vectors. Moreover, by avoiding explicit CE at the UE, JEFPNet reduces both storage, distribution, and computational overhead.
@@
Notations:In this paper, we denote Vectors and Matrices with boldface lower- and upper-case letters, respectively. ${\mathbf{A}}^{T}$ and ${\mathbf{A}}^{H}$ are transpose of $\mathbf{A}$} and complex conjugate transpose operation of $\mathbf{A}$}, respectively. $a^{*}$ denotes the conjugate processing of complex variable $a$. $\left\|\cdot\right\|_{1}$}, $\left\|\cdot\right\|_{2}$} and $\lVert\cdot\rVert$ represents the -$\ell\text{-}1$ norm, the -$\ell\text{-}2$ norm and modulus operations on complex numbers, respectively. $\mathbb{E}\left(\cdot\right)$ denotes the statistical expectation. $\{\cdot,\cdot\}$ denotes a set containing multiple elements. $\{\mathbf{A}_{k}\}_{k=1}^{K}$ denotes a set containing $K$ elements from $\mathbf{A}_{1}$ to $\mathbf{A}_{K}$, and $\{\mathbf{A}(k)\}_{k=1}^{K}$ denotes a set containing $K$ elements from $\mathbf{A}(1)$ to $\mathbf{A}(K)$.The symbols $\mathbb{R}$ and $\mathbb{C}$ represent sets of real and complex numbers, respectively. $\lfloor{A}\rfloor$ indicates the floor operation.
@@ II-A System Model
Given that OFDM systems underpin most contemporary broadband wireless standards (e.g., long term evolution (LTE), 5G new radio (NR) ) and offer well-established advantages in multiuser MIMO scenarios, we adopt a MIMO–OFDM system with $K$ users and a maximum of $K_{\max}$ users for our proposed scheme to ensure practical deployability. Each user is equipped with a single omnidirectional antenna, while the BS employs an $N_{t}$-element antenna array. Although one could first investigate a narrowband system and subsequently extend it to a wideband OFDM system through a parallel scheme, such an approach fails to exploit frequency-domain correlation and introduces additional learning parameters. Therefore, we directly study the problem of multiuser precoding in a wideband system. The system operates with $N_{c}$ subcarriers. In this system, we simultaneously consider downlink CE, downlink CSI feedback in the uplink channel, and downlink precoding design. The downlink CSI of the $k$-th user can be expressed as $\mathbf{H}_{d}(k)\in\mathbb{C}^{N_{c}\times N_{t}}$, and the corresponding uplink feedback channel can be expressed as $\mathbf{H}_{u}(k)\in\mathbb{C}^{N_{c}\times N_{t}}$.
For the downlink CE phase, we consider the pilot-assisted CE method, that downlink CSI is estimated at the UE with the help of the known pilot sequence for transmitter and receiver. For the pilot sequence, in the time domain, we use sequential $L$ OFDM symbols for CE. To reduce pilot overhead, pilots are placed at equal $g$ intervals in the frequency domain, while the non-pilot positions are utilized for data or other channel control information transmission. $M$ subcarriers are sampled at equal intervals $g$ from the $N_{c}$ subcarriers to place the pilots, and the indices of these subcarriers form the set of pilot sequences, denoted as $\mathcal{M}=\{\mathcal{M}_{1},\cdots,\mathcal{M}_{M}\}\in\mathbb{R}$, where $M=\lfloor{N_{c}/g}\rfloor$. In this paper, user interference during feedback is not considered. Consequently, the same pilot symbols are employed across different users. The pilot sequence at the $\mathcal{M}_{m}$-th subcarrier can be expressed as $\mathbf{P}_{\mathcal{M}_{m}}=\left[{\mathbf{p}_{\mathcal{M}_{m}}^{(1)},\cdots,\mathbf{p}_{\mathcal{M}_{m}}^{(L)}}\right]\in\mathbb{C}^{N_{t}\times L}$.
@@
At the BS, we use a maximal ratio combining (MRC) receiver for detecting signals. The combining matrix is derived by the estimated uplink CSI. In this paper, we assume perfect uplink CE. Therefore, the combining matrix of the $z$-th subcarrier $\mathbf{W}_{z}(k)$ can be expressed as
@@
It is worth emphasizing that, for simplicity, we considers a fully digital beamforming architecture in this paper, where the number of radio frequency (RF) chains $N_{\mathrm{RF}}$ is equal to the number of BS antennas $N_{t}$. If the proposed network is to be applied in a hybrid beamforming architecture, the analog precoder $\mathbf{F}_{\mathrm{RF}}\in\mathbb{C}^{N_{t}\times N_{\mathrm{RF}}}$ and the digital precoder $\mathbf{F}_{\mathrm{BB}}\in\mathbb{C}^{N_{\mathrm{RF}}\times K}$ should be designed separately for effective deployment.